# Redistribution Through

### v2

@@ Abstract
We study how to optimally segment a monopolistic market given a redistributive objective. Optimal redistributive segmentations (i) induce the seller to price progressively, i.e., richer consumers pay higher prices than poorer ones, and (ii) may require giving a higher profit than uniform pricing if the redistributive motive is strong. We further show that optimal redistributive segmentations are implementable via price-based regulation.
@@ 1 Introduction
Firms and online platforms use rich consumer data to segment markets and personalize prices. While these practices can improve economic efficiency and even increase total consumer surplus, questions remain about their distributional consequences. Since price discrimination affects consumers differently, assessing its welfare effects through aggregate consumer surplus conceals its heterogeneous impacts. In fact, segmenting markets so as to maximize total consumer surplus may disproportionately benefit wealthier consumers.
@@
As Bergemann et al. (2015) show, monopolistic markets can be segmented so that the whole generated surplus accrues to consumers. This outcome is achieved by creating market segments that pool consumers with high and low willingness to pay. These segmentations exploit pooling externalities (Galperti et al., 2023): consumers with low willingness to pay exert a positive externality on consumers with high willingness to pay in the same segment by inducing the seller to charge prices below the uniform monopoly price. As a byproduct, buyers with a high willingness to pay capture most of the surplus. Nevertheless, a low willingness to pay might reflect adverse economic and social circumstances. Consumer surplus maximization may, therefore, unintentionally benefit wealthy buyers to the detriment of poorer ones.
@@
In this paper, we study optimal segmentations for an objective that prioritizes the welfare of consumers with low willingness to pay. Our two main economic insights are the following. First, unlike segmentations that maximize total consumer surplus—which assign consumers with high willingness to pay to low-price segments—, optimal redistributive segmentations are positive-assortative: they assign buyers with high willingness to pay to high-price segments and buyers with low willingness to pay to low-price segments. As a result, they induce the monopolist to price progressively, meaning that wealthier consumers pay higher prices than poorer ones (Section 3.4). Second, this separation of wealthy and poor consumers into different segments might allow the seller to price discriminate more effectively and increase its profit compared to uniform pricing. Therefore, redistributive segmentations may inadvertently allocate part of the surplus created to the seller, thereby not maximizing total consumer surplus (Section 3.5).
Our model considers a continuous population of buyers, each having a unit demand for a good sold by a monopolistic seller. Each buyer is characterized by a type equal to their willingness to pay for the good. We refer to the initial distribution of types in the economy as the aggregate market. [^1] We follow the terminology of Bergemann et al. (2015). Without loss of generality, we assume that the seller’s cost of production is zero, so gains from trade are always positive. A market designer can flexibly segment the aggregate market, anticipating that the seller will charge a profit-maximizing price in each segment. The designer’s preferences are represented by a welfare function, which describes the social value of each buyer type consuming at each possible price. We restrict attention to welfare functions that align with consumer surplus: welfare is positive if and only if the buyer is willing to consume, and welfare is decreasing in price. Among those, we consider redistributive welfare functions: any price discount is more welfare-improving when granted to a low-type buyer rather than a high-type buyer. Mathematically, this property corresponds to supermodularity. Two examples of redistributive welfare functions are the weighted utilitarian welfare function with decreasing Pareto weights, which have received particular attention in the recent literature on redistributive mechanism design (Dworczak et al., 2021; Akbarpour ⓡ al., 2024b); and concave increasing transformations of consumer surplus, considered in the literature on optimal taxation (Mirrlees, 1971) and insurance (Rothschild and Stiglitz, 1976).
We begin by showing that optimal redistributive segmentations must be efficient (Lemma 1), so that they induce the seller to serve all buyers. Therefore, there is no equity-efficiency trade-off. [^2] This result is implied by the fact that any inefficient segmentation can be Pareto-improved (Haghpanah and Siegel, 2023). We then define the redistributive order, a partial order that ranks efficient segmentations according to their degree of redistributivity. Specifically, a segmentation is more redistributive than another if it yields higher aggregate welfare under any redistributive welfare function. Our first main result, Theorem 1, characterizes this order in terms of buyer mass transfers between market segments. We show that a segmentation is more redistributive than another if and only if it can be derived from the latter through a finite sequence of two kinds of transfers: (i) downward transfers, that move a certain mass of consumers to a lower-price segment; (ii) redistributive transfers, that swap a mass of buyers of a given type in a high-price segment with an equal mass of higher-type buyers in a lower-price segment. Downward transfers increase welfare because redistributive welfare functions decrease in price, whereas redistributive transfers increase welfare because redistributive welfare functions satisfy supermodularity. [^3] Our characterization of the redistributive order in terms of mass transfers is analogous to the characterization of the supermodular stochastic order carried out independently by Müller (2013) and Meyer and Strulovici (2015). For further discussion, see Section 3.2.
We then prove in Proposition 1 that the set of optimal redistributive segmentations is exactly the set of segmentations that are undominated in the redistributive order. Specifically, our definition of the redistributive order implies that, for any redistributive welfare function, an undominated optimal segmentation exists. Proposition 1 shows the converse statement: any undominated segmentation can be rationalized as optimal for some redistributive welfare function. Building on this equivalence, Proposition 2 further characterizes optimal redistributive segmentations. Theorem 1 implies that undominated segmentations are such that no downward or redistributive transfer can be performed without inducing the seller to charge a higher price in at least one segment. Proposition 2 identifies the set of prices between which the seller must be indifferent, so that performing any downward or redistributive transfer violates at least one of the seller’s incentive constraints.
@@
We show the optimality of progressive pricing by proving that any optimal segmentation must assign buyer types to prices in a positive-assortative manner. Specifically, Theorem 2.A establishes that any optimal segmentation satisfies a weak form of monotonicity, which requires that, if a buyer consumes at a given price, there must exist a buyer with a greater willingness to pay who consumes at a higher price. Optimal segmentations may also satisfy a stronger notion of monotonicity: if a buyer consumes at a given price, then all buyers with a greater willingness to pay consume at a higher price. We show in Theorem 2.B that optimal segmentations satisfy this strong form of monotonicity if and only if the welfare function satisfies a strengthening of supermodularity reflecting strong redistributive concerns. Furthermore, optimal strongly monotone segmentations are obtained by following a simple greedy algorithm that iteratively maximizes the welfare of every type starting from the lowest.
@@
Next, we show that redistributive segmentations may not maximize consumer surplus. The monotonicity of redistributive segmentations separates buyers with low and high willingness to pay into different segments, possibly allowing the seller to price-discriminate more effectively than under consumer-optimal segmentations. As a result, the seller might obtain a strictly positive share of the surplus generated by segmenting the market. In such cases, we say that the seller enjoys a redistributive rent. Redistributive rents arise especially when optimal segmentations satisfy the strong form of monotonicity described in the previous paragraph. Theorem 3 characterizes the markets for which strong redistributive preferences necessarily induce a redistributive rent for the seller. When rents are not necessary, the optimal segmentation is very simple with only two segments: one where the lowest possible price is charged, pooling all types that would not have bought the good in the unsegmented market; and one complementary segment where the uniform monopoly price is charged.
@@
Our analysis considers the segmentation problem faced by a market designer with redistributive concerns. In practice, however, market segmentation is primarily conducted by private actors whose objectives are not necessarily aligned with welfare maximization. For instance, online platforms may focus on maximizing the number of transactions (Romanyuk and Smolin, 2019), while data intermediaries aim for profit maximization (Yang, 2022). In the last section of the paper, we hence investigate under what conditions redistributive segmentations can be implemented by a regulator who can only imperfectly observe the segmentation chosen by an agent, assumed here to be the seller. [^4] The result would still hold if the agent had a different objective function, provided it is aligned with efficiency. For example, a platform remunerated by a flat fee on transactions would seek to maximize participation and thus segment efficiently. We introduce a notion of price-based implementation: a segmentation is implementable through price-based regulation if there exists no alternative segmentation that strictly increases the seller’s profit while inducing the same distribution of prices. [^5] This notion of implementation is equivalent to Lin and Liu’s (2024) notion of credibility. We further discuss the connection to that contribution in Section 4, Footnote 28. This notion of implementation is motivated by the fact that although regulators are likely to possess less information than data intermediaries, they can plausibly observe the distribution of charged prices. [^6] Many websites, such as Camelcamelcamel, Keepa, Price Before, and Honey, already specialize in tracking the evolution, frequency and extent of price fluctuations for products sold on Amazon. In the context of regulating online platforms, for instance, it only involves monitoring transaction data and imposing penalties based on the divergence of the empirical price distribution from the desired target distribution. In Proposition 3, we show that any optimal redistributive segmentation is implementable through price-based regulation. [^7] In fact, Proposition 3 states a more general result of independent interest: any efficient segmentation can be implemented via price-based regulation.
@@
Our paper contributes to the analysis of the welfare effects of third-degree price discrimination, following Pigou (1920) and Robinson (1933). Much of this literature focuses on studying how given market segmentations affect efficiency, consumer surplus, total output and prices (Schmalensee, 1981; Varian, 1985; Aguirre et al., 2010; Cowan, 2016; Rhodes and Zhou, 2024). Le Grand (1975) is the first to study theoretically the use of third-degree price discrimination as a redistributive device for public firms.
@@
Bergemann et al. (2015) pioneered an information-design-based approach to third-degree price discrimination by characterizing the feasible welfare outcomes under any segmentation of a monopolistic market. [^8] We refer to Kamenica (2019) and Bergemann and Morris (2019) for comprehensive reviews of the Bayesian persuasion and information design literature. They show that any combination of consumer surplus and producer profit such that the seller’s profit is at least equal to that under uniform pricing, the consumer surplus is non-negative, and the total surplus does not exceed the efficient level, is achievable through some segmentation. Haghpanah and Siegel (2022) and Terstiege and Vigier (2023) perform the same exercise in a multi-product setting, while Elliott et al. (2024) and Bergemann et al. (2024) consider competitive markets. Kartik and Zhong (2023), extend the characterization of welfare outcomes of Bergemann et al. to the case where the buyer and the seller have correlated values. [^9] A parallel literature following Roesler and Szentes (2017) assumes that buyers are imperfectly informed about their valuation and characterize the consumer-optimal information structures. Condorelli and Szentes (2020) characterize consumer-optimal information structures in the hold-up problem. Ravid et al. (2022) consider the case where the buyer has a cost of acquiring information. Deb and Roesler (2021) characterize consumer-optimal information structures in the multidimensional screening problem. Armstrong and Zhou (2022) characterize consumer-optimal information structures when the downstream market is a differentiated product duopoly. We contribute to this literature by studying the impact of market segmentation on consumer welfare beyond the standard of consumer surplus. In particular, we study redistributive welfare functions and show that prioritizing low-type consumers does not hinder efficiency but may not maximize total consumer surplus.
Doval and Smolin (2024) study the welfare effects of information provision on a heterogeneous population of agents. They introduce the Bayes welfare set, consisting of the set of welfare profiles that can be achieved under any information policy. They characterize this set geometrically and show that its Pareto frontier is generated by the set of solutions of a family of Bayesian persuasion problems, in which a weighted utilitarian social planner plays the role of the designer. As such, they characterize the welfare effects that any segmentation can have on consumers of different types, without assuming any particular shape for the welfare function. In our paper, we assume the welfare function satisfies particular conditions—specifically, redistributivity—, solve explicitly for the optimal segmentations and show that their structure differs from Bergemann et al.’s consumer-surplus maximizing segmentations.
Dubé and Misra (2023) conduct a randomized controlled field experiment to study the welfare implications of personalized pricing. Although they find a negative effect of personalized pricing on total consumer surplus, the effect on aggregate consumer welfare is positive under some inequality-averse welfare functions. This shows that low-valuation buyers are the ones who benefit from the implemented price discrimination. Buchholz et al. (2020) study the effect of price discrimination on a two-sided platform for cab rides. They also find that while aggregate consumer surplus decreases, more price-sensitive buyers benefit from personalized pricing.
@@
Following the seminal contribution of Weitzman (1977), a burgeoning literature has integrated redistributive concerns into mechanism design (see, e.g., Condorelli, 2013; Reuter and Groh, 2020; Dworczak et al., 2021; Pai and Strack, 2022; Akbarpour ⓡ al., 2024a, b; Kang, 2023; Kang and Watt, 2024). [^10] Dworczak (2024) gives an overview of the recent contributions to inequality-aware mechanism design. Our paper studies redistribution in an information design setting. [^11] Arya and Malhotra (2022) studies redistribution through a different notion of market segmentation, namely, the non-tradability of essential and non-essential goods in a general equilibrium framework. Contrary to the optimal mechanisms in this literature, our redistributive segmentations are socially efficient and thus do not exhibit any equity-efficiency trade-off.
2 Model
Economy
@@ 2 Model
A monopolistic seller (she) supplies a good to a unit mass of buyers. Buyers have a unit demand and differ in their willingness to pay for the good, $\theta$, assumed to belong to the finite set $\Theta=\{\theta_{1},\dots,\theta_{K}\}\subset\mathbb{R}_{+}$. We refer to the willingness to pay as the buyer’s type and use the two terms interchangeably. A buyer willing to pay $\theta$ consumes at a price $p\in\mathbb{R}_{+}$ if and only if $\theta\geq p$. We assume that the seller’s marginal cost of production is zero, so gains from trade are positive. [^12] This assumption is without loss of generality. If the seller’s marginal cost were equal to $c>0$, then one could redefine the buyer type as $\tilde{\theta}=\max\{0,\theta-c\}$ for all $\theta\in\Theta$. The aggregate market (in short the market) $\mu\in\Delta(\Theta)$ specifies the initial distribution of types.
@@
Condition $\mathrm{C}_{\mu}$ ensures that the mass of type $\theta$ buyers assigned to all segments coincides with the mass of those types in the market. [^13] We can equivalently think of any pair $(\sigma,S)$ as a Blackwell experiment which, for any prior distribution $\mu$ over the state space $\Theta$, describes the joint distribution of signals $s\in S$ and states of the world $\theta\in\Theta$, and satisfies the Bayes-plausibility constraint $\mathrm{C}_{\mu}$ (Aumann et al., 1995; Kamenica and Gentzkow, 2011). According to that interpretation, each $\sigma$ represents a way of revealing information about the buyers’ types to the seller.
@@
When facing a market segmentation, the seller sets prices so as to maximize his profit on each generated segment. That is, for any $\mu\in\Delta(\Theta)$, any $\sigma\in\Sigma(\mu)$, and any $s\in S$ such that $\sigma(\theta,s)>0$ for some $\theta\in\Theta$, the seller solves
@@
where the set of prices is given by $P=\Theta$. Indeed, there is no loss of generality in assuming that the seller charges prices in $\Theta$ since, for any $k\in\{1,\dots,K\}$, charging a price falling strictly between two consecutive types $\theta_{k}$ and $\theta_{k+1}$ always yields a lower profit to the seller than charging either $p=\theta_{k}$ or [^14] It is also easy to see that charging any price lower than $\theta_{1}$ or greater than $\theta_{K}$ is strictly dominated. $p=\theta_{k+1}$.
Direct segmentations
Following Proposition 2 in Bergemann et al. (2015), it is without loss of generality to focus on direct market segmentations. A segmentation is direct if each price is charged in at most one segment. Direct segmentations correspond to direct communication mechanisms in the sense of Myerson (1982). In such mechanisms $S=P$, so that messages correspond to price recommendations to the seller.
A direct segmentation is obedient if the seller finds it optimal to follow the price recommendation in each segment. That is:
@@
For any $\mu\in\Delta(\Theta)$, we denote the set of direct and obedient segmentations of $\mu$ as $\Sigma(\mu)$. We henceforth slightly abuse terminology by referring to any element of $\Sigma(\mu)$ as a segmentation of $\mu$. For any $\mu\in\Delta(\Theta)$ and any $\sigma\in\Sigma(\mu)$, we let $\sigma_{P}\in\Delta(P)$ be the price distribution induced by $\sigma$, defined by:
@@
Let $\mu\in\Delta(\Theta)$. An instance of direct segmentation of $\mu$ is illustrated on Figure 1.
$\theta$$p$$(\theta_{1},\theta_{1})$$p^{\star}_{\mu}$$\theta_{n}$$p^{\star}_{\mu}$$\theta_{n}$$p^{\prime}$$p^{\prime\prime}$
@@
The initial mass of consumers, which is located at the row $p^{\star}_{\mu}$ in the grid $\Theta\times P$, is split between two segments $p^{\prime}$ and $p^{\prime\prime}$ in a vertical way, so that $\mathrm{C}_{\mu}$ is satisfied. Furthermore, Ob requires that the mass moved at each point must be such that $p^{\prime}$ and $p^{\prime\prime}$ are respectively profit-maximizing prices in segments $p^{\prime}$ and $p^{\prime\prime}$.
@@
We consider a designer (he) who can segment the market flexibly. His preferences over market outcomes are represented by a welfare function $w\colon\Theta\times P\to\mathbb{R}$. For any welfare function $w$ and segmentation $\sigma\in\Sigma(\mu)$, the aggregate welfare is

### v5

@@ Abstract
We study how to optimally segment monopolistic markets with a redistributive objective. We characterize optimal redistributive segmentations and show that they (i) induce the seller to price progressively, i.e., charge richer consumers higher prices than poorer ones, and (ii) may not maximize consumer surplus, instead granting extra profits to the monopolist. We further show that optimal redistributive segmentations are implementable via price-based regulation.
@@ 1. Introduction
Firms and online platforms use rich consumer data to segment markets and personalize prices. [^1] Third-degree price discrimination is often implemented through targeted discounts. For instance, this is how Amazon Brand Tailored promotions and Google promotion assets operate. See mohammed17how; wallheimer18are; oecd2022personalised for more details and evidence. While these practices can improve economic efficiency and even increase total consumer surplus, questions remain about their distributional consequences.
Distributive concerns have long shaped both policy practice and fairness norms regarding price discrimination. From “lifeline” tariffs to student discounts at cultural activities or food stamps to low-income households, price discrimination has repeatedly been justified on the basis of who it benefits. Recent regulatory debates on online personalized pricing have expressed similar concerns (WhiteHouse2015; bourreau2018regulation). Despite this, most of the theoretical literature on price discrimination has focused on studying its impacts on aggregate measures such as output or total consumer surplus. Price discrimination inherently affects consumers differently, and evaluating its welfare effects solely through such aggregates conceals these heterogeneous impacts.
@@
In particular, segmenting markets to maximize total consumer surplus can favor wealthier consumers. As bbm15 show, monopolistic markets can be segmented so that the whole generated surplus accrues to consumers. This outcome is achieved by exploiting pooling externalities (galperti2023): consumers with a low willingness to pay exert a positive externality on consumers with a high willingness to pay who belong to the same segment by inducing the seller to charge prices below the uniform monopoly price. Adequately pooling consumers in the same segments therefore increases the total share of the surplus going to consumers, particularly those most willing to pay. However, a low willingness to pay is often indicative of adverse economic and social circumstances. Therefore, maximizing consumer surplus can benefit wealthy buyers to the detriment of poorer ones.
@@
This paper aims to better understand these distributive effects of market segmentation. Specifically, we address the normative question of how to segment a market to benefit consumers while prioritizing the welfare of those with low willingness to pay. Our two main economic insights are the following. First, unlike segmentations that maximize total consumer surplus, which assign consumers with a high willingness to pay to low-price segments, optimal redistributive segmentations are positive-assortative: they assign buyers with a high willingness to pay to high-price segments and buyers with low willingness to pay to low-price segments. As a result, they induce the monopolist to price progressively, meaning that wealthier consumers pay higher prices than poorer ones (Section˜3.3). Second, this separation of wealthy and poor consumers in different segments sometimes allows the seller to price discriminate more effectively, and therefore increase its profit compared to uniform pricing. Hence, redistributive segmentations may not maximize total consumer surplus (Section˜3.4).
Our model considers a continuous population of buyers, each having a unit demand for a good sold by a monopolistic seller. Every buyer is characterized by her willingness to pay for the good—her type. We refer to the distribution of types in the economy as the aggregate market. Without loss of generality, we assume that the seller’s cost of production is zero, so gains from trade are always positive. A market designer can flexibly segment the aggregate market, anticipating that the seller will charge a profit-maximizing price in each segment.
The designer’s preferences are represented by a welfare function, which describes the social value of a given type consuming at a given price. We focus on welfare functions that align with consumer surplus: welfare equals zero when buyers do not purchase, and is decreasing in price when buyers purchase. Within this class, we study redistributive welfare functions, for which the welfare gain resulting from reducing the price paid by buyers with low type is greater than the welfare gain resulting from an equal price reduction for buyers with high type. [^2] Mathematically, this corresponds to the welfare function being supermodular. This is motivated by the idea that buyers’ willingness to pay reflects both their taste for the good and their opportunity cost of money. Prioritizing low willingness to pay buyers thus appropriately captures redistributive concerns when heterogeneity in willingness to pay primarily reflects differences in ability to pay: wealthier individuals can typically pay more for what they need. [^3] We refer to dworczak21; akbarpour2024redistributive; dworczak2024inequality for in-depth discussions of this argument. Examples of redistributive welfare functions include weighted consumer surplus with Pareto weights decreasing in willingness to pay, or increasing concave transformations of consumer surplus.
Our main result, Theorem˜1, characterizes the set of segmentations that are optimal for some redistributive welfare function as those that saturate a specific set of the seller’s incentive constraints. These binding incentive constraints ensure that no marginal reallocations of buyers across segments can be beneficial for every redistributive welfare function. Theorem˜1 relies on a key result of independent interest. We define a redistributive order, which ranks a segmentation as more redistributive than another if it yields higher aggregate welfare for every redistributive welfare function. We characterize it in terms of elementary reallocations of buyers across segments: a segmentation is more redistributive than another if and only if there exists a sequence of such elementary reallocations transforming the latter into the former (Proposition˜1).
@@
We leverage Theorem˜1 to draw key economic implications of optimal redistributive segmentations. We first show the optimality of progressive pricing by proving that any optimal segmentation must assign buyer types to prices in a positive-assortative manner. Specifically, Corollary˜1 establishes that any optimal segmentation satisfies a weak form of monotonicity, which requires that, if a buyer consumes at a given price, there must exist a buyer with a greater willingness to pay who consumes at a higher price. This notion of progressive pricing is mild, since it still allows for instances in which some buyer with higher willingness to pay consumes at a price lower than some buyer with lower willingness to pay. We show in Proposition˜2 how a strengthening of redistributive concerns [^4] As shown below, this corresponds mathematically to the welfare function satisfying a strengthened notion of supermodularity (Definition 6). eliminates these instances, leading to a stronger notion of monotonicity: if a buyer consumes at a given price, then all buyers with a (weakly) greater willingness to pay consume at a higher price. Furthermore, we show that optimal strongly monotone segmentations are obtained by following a simple greedy algorithm that iteratively maximizes the welfare of every type starting from the lowest.
@@
Second, while redistribution does not induce an equity-efficiency trade-off (Lemma˜1 shows that optimal redistributive segmentations are efficient), we show that it may be in conflict with consumer surplus maximization. The monotonic structure of redistributive segmentations implies separation of buyers with low and high willingness to pay into different segments, possibly allowing the seller to price discriminate more effectively than under consumer surplus maximizing segmentations. As a result, the seller might obtain a strictly positive share of the surplus generated by segmenting the market. In such cases, we say that the seller enjoys a redistributive rent. Redistributive rents arise especially when optimal segmentations satisfy the strong form of monotonicity described in the previous paragraph. Proposition˜3 characterizes the markets for which strongly redistributive preferences necessarily induce a redistributive rent for the seller. When rents are not necessary, the optimal segmentation is remarkably simple. It only generates two segments: one where the lowest possible price is charged, pooling all types that would not have bought the good in the unsegmented market; and one segment where the uniform monopoly price is charged.
@@
Our analysis considers the problem faced by a designer with redistributive concerns who can flexibly segment the aggregate market. In practice, however, market segmentation is primarily conducted by private actors whose objectives are not necessarily aligned with welfare maximization. For example, online platforms can focus on maximizing the number of transactions (Romanyuk2019), while data intermediaries aim to maximize profits (Yang2022). In the last section of the paper, we therefore investigate under what conditions redistributive segmentations can be implemented by a regulator who can only imperfectly observe the segmentation chosen by an agent, assumed here to be the seller. [^5] The result would still hold if the agent had a different objective function, provided it is aligned with efficiency. For example, a platform remunerated by a flat fee on transactions would seek to maximize participation, and thus segment efficiently. We introduce a notion of price-based implementation: a segmentation is implementable through price-based regulation if there exists no alternative segmentation that strictly increases the seller’s profit while inducing the same distribution of prices. [^6] This notion of implementation is equivalent to linliu’s (linliu) notion of credibility. We further discuss the connection to that contribution in Section 4, Footnote 35. This notion of implementation is motivated by the fact that although regulators are likely to possess less information than data intermediaries, they can plausibly observe the distribution of charged prices. [^7] Many websites, such as Camelcamelcamel, Keepa, Price Before, and Honey, already specialize in tracking the evolution, frequency and extent of price fluctuations for products sold on Amazon. In the context of regulating online platforms, for instance, it only involves monitoring transaction data and imposing penalties based on the divergence of the empirical price distribution from the desired target distribution. In Proposition˜4, we show that any optimal redistributive segmentation is implementable through price-based regulation. [^8] In fact, Proposition 4 states a more general result of independent interest: any efficient segmentation can be implemented via price-based regulation.
@@
Our paper contributes to the analysis of the welfare effects of third-degree price discrimination, following pigou and robinson. Much of this literature focuses on studying how given market segmentations affect efficiency, consumer surplus, total output, and prices (schmalensee; varian; Aguirre2010Monopoly; Cowan2016Welfare; rhodeszhou).
legrand is the first to study theoretically the use of third-degree price discrimination as a redistributive device for public firms. His analysis is expanded by steinberg to the case where there is competition between for-profit firms and a (distributionally minded) non-profit provider.
@@
bbm15 pioneered an information design approach to third-degree price discrimination. [^9] See kamenicareview and Bergemann2019 for comprehensive reviews of the Bayesian persuasion and information design literature. They show that any pair of consumer surplus and producer profit is attainable through some segmentation if the seller’s profit is at least the uniform-pricing level, consumer surplus is non-negative, and total surplus does not exceed the efficient level. hagpanah_siegel_aer and terstiege2023 extend this analysis to multi-product environments, elliott2024Market and bergemann2024alignment study competitive markets, and Salas2025 examines how optimal segmentation is shaped by endogenous buyer participation. We contribute to this literature by studying the impact of market segmentation on consumer welfare beyond the standard of consumer surplus. In particular, we study redistributive welfare functions and show that prioritizing low-type consumers does not hinder efficiency but may not maximize total consumer surplus.
@@
A burgeoning literature rooted in the seminal contribution of weitzman1977 shows how markets (dworczak21), allocation mechanisms (Condorelli2013; Reuter2020; akbarpour2024vaccines; akbarpour2024redistributive; kang2023; kang2024a; kang2024b), or commodity taxation (pai2022taxing; Ahlvik2025; Doligalski2025) can be optimally designed to mitigate economic inequality.101010dworczak2024inequality provides an overview of recent contributions to inequality-aware market design. Our paper contributes to this literature by showing how market segmentation (or, equivalently, information) can serve as a redistributive device.111111arya2022rethinking study redistribution through a different notion of market segmentation, namely the non-tradability of essential and non-essential goods in a general equilibrium framework.
Relatedly, doval_smolin_22 analyze how information provision affects the welfare of heterogeneous agents. They introduce the Bayes welfare set, defined as the set of agents’ welfare profiles that can be achieved under some information policy. Their framework applies to a broad class of persuasion environments, including the monopolistic market we study. They characterize this set geometrically and show that its Pareto frontier is generated by the solutions to a family of Bayesian persuasion problems in which a weighted-utilitarian social planner plays the role of the designer. This delivers a complete description of the welfare effects that any segmentation can have on consumers of different types.
Rather than characterizing which welfare profiles are feasible, we study how to optimally distribute consumer welfare through information design. To do so, we specify a class of redistributive welfare functions—those that are supermodular in the types and the prices—and fully characterize all segmentations that are optimal for such objectives, for any distribution of consumer types. This characterization implies the use of monotone information structures as optimal policies—structures that are often viewed as especially attractive in the persuasion literature because of their simplicity (see, e.g., Ivanov2021; Mensch2021; Nikzad2021; Kolotilin2022; Onuchic2023; Kolotilin2024). This literature often obtains monotone structures as optimal in the case where payoffs are linear in the state (that is, only the posterior mean matters). [^12] Notable exceptions are, e.g., Mensch2021 and Nikzad2021. This is not the case in the environment we consider—both the sender’s and receiver’s payoffs are non-linear. We thus develop a new method inspired by the literature on stochastic orders (cf. Proposition˜1) to prove our characterization. [^13] We discuss the connection between our approach and the literature on stochastic orders and optimal transport in greater detail in Section 3.5.2.
The distributive impacts of price discrimination have also been investigated empirically. Dube2023 conduct a randomized controlled field experiment to study the welfare implications of personalized pricing. While they find that personalized pricing reduces total consumer surplus, aggregate consumer welfare can increase under inequality-averse welfare criteria, indicating that buyers with low willingness to pay benefit the most from the implemented price discrimination. Similarly, buchholz2020value examine price discrimination on a two-sided platform for cab rides and find that, although aggregate consumer surplus falls, more price-sensitive buyers gain from personalized pricing.
2. Model
Primitives
@@ 2. Model
A monopolistic seller (she) supplies a good to a unit mass of buyers. Buyers have a unit demand and differ in their willingness to pay for the good, $\theta$, assumed to belong to the finite set $\Theta=\{\theta_{1},\dots,\theta_{K}\}\subset\mathbb{R}_{+}$, where $\theta_{1}<\dots<\theta_{K}$. We refer to the willingness to pay as the buyer’s type and use the two terms interchangeably. A buyer willing to pay $\theta$ consumes at a price $p\in\mathbb{R}_{+}$ if and only if $\theta\geq p$. We assume that the seller’s marginal cost of production is zero, so the gains from trade are positive. [^14] This assumption is without loss of generality. If the seller’s marginal cost were equal to $c>0$, then one could redefine the buyer type as $\tilde{\theta}=\max\{0,\theta-c\}$ for all $\theta\in\Theta$. The aggregate market (in short the market) $\mu\in\Delta(\Theta)$ specifies the initial distribution of types.
@@
Condition $\mathrm{C}_{\mu}$ ensures that the mass of type $\theta$ buyers assigned to all segments coincides with the mass of those types in the market. [^15] We can equivalently think of any pair $(\sigma,S)$ as a Blackwell experiment which, for any prior distribution $\mu$ over the state space $\Theta$, describes the joint distribution of signals $s\in S$ and states of the world $\theta\in\Theta$, and satisfies the Bayes-plausibility constraint $\mathrm{C}_{\mu}$ (Aumann1995; kamenica_gentzkow11). According to that interpretation, each $\sigma$ represents a way of revealing information about the buyers’ types to the seller.
@@
When faced with a market segmentation, the seller sets prices to maximize her profit in each segment. That is, for any $\mu\in\Delta(\Theta)$, any $\sigma$ satisfying $\mathrm{C}_{\mu}$, and any $s\in S$ such that $\sigma(\theta,s)>0$ for some $\theta\in\Theta$, the seller solves
@@
where the set of prices is given by $P=\Theta$. Indeed, there is no loss of generality in assuming that the seller charges prices in $\Theta$ since, for any $k\in\{1,\dots,K\}$, charging a price falling strictly between two consecutive types $\theta_{k}$ and $\theta_{k+1}$ always yields a lower profit to the seller than charging either $p=\theta_{k}$ or $p=\theta_{k+1}$. [^16] Similarly, charging any price lower than $\theta_{1}$ or greater than $\theta_{K}$ is never optimal.
Direct and obedient segmentations
It is without loss of generality to focus on direct and obedient market segmentations. [^17] Precisely, it is without loss given the objectives we define below. See Bergemann2019. In such segmentations, $S=P$, so that the segments are indexed by a recommended price, and the seller finds it optimal to follow this recommendation. That is:
@@
For any $\mu\in\Delta(\Theta)$, we denote the set of direct and obedient segmentations of $\mu$ as $\Sigma(\mu)$. We henceforth slightly abuse terminology by referring to any element of $\Sigma(\mu)$ as a segmentation of $\mu$. For any $\mu\in\Delta(\Theta)$ and any $\sigma\in\Sigma(\mu)$, we let $\sigma_{P}\in\Delta(P)$ be the distribution of recommended prices induced by $\sigma$, defined by: [^18] As will become clear after Lemma 1 is presented, in optimal segmentations the distribution of recommended prices will coincide with the distribution of realized transaction prices.
@@
Let $\mu\in\Delta(\Theta)$. An example of direct segmentation of $\mu$ is illustrated in Figure˜1.
[…]
@@
The initial mass of consumers, located in the row $p^{\star}_{\mu}$ in the grid $\Theta\times P$, is divided between two segments $p^{\prime}$ and $p^{\prime\prime}$ vertically, so $\mathrm{C}_{\mu}$ is satisfied. Furthermore, Ob requires that the mass moved at each point be such that $p^{\prime}$ and $p^{\prime\prime}$ are, respectively, profit-maximizing prices in segments $p^{\prime}$ and $p^{\prime\prime}$.
@@
We consider a designer (he) who can segment the market flexibly. His preferences over market outcomes are represented by a welfare function $w\colon\Theta\times P\to\mathbb{R}$. For any welfare function $w$ and segmentation $\sigma\in\Sigma(\mu)$, the aggregate welfare is defined by